solve by method of undetermined coefficient : y'' + 2(y') + 5y = x(e^-x)cos(2x)
1
Expert's answer
2020-03-25T12:39:57-0400

"y''+2y'+5y=xe^{-x}\cos(2x)"

General solution will be sum of the complementary solution and particular solution.

Find the complementary solution by solving "y''+2y'+5y=0"

Assume a solution will be proportional to "e^{\lambda x}" for some constant "\lambda"

Substitute "y(x)=e^{\lambda x}" into the differential equation

"(\lambda^2+2\lambda+5)e^{\lambda x}=0"

solve for "\lambda" : "\lambda = -1+2i" or "\lambda=-1-2i"

These roots give this general solution

"y_g(x)=c_1e^{-x}\cos{2x}+c_2e^{-x}\sin{2x}"

To find partial solution, we put into equation this form of solution

"y_p(x)=axe^{-x}\cos(2x)+bxe^{-x}\sin(2x)+cx^2e^{-x}\cos(2x)+dx^2e^{-x}\sin(2x)"

and by computing first and second derivative and putting everything into equation, we determine all the coefficients

Partial solution:

"y_p(x)=1/16xe^{-x}\cos{2x}+1/8x^2e^{-x}\sin{2x}"

Then "y(x)=y_g(x)+y_p(x)=c_1e^{-x}\cos{2x}+c_2e^{-x}\sin{2x}+1/16xe^{-x}\cos{2x}+1/8x^2e^{-x}\sin{2x}"